### v1

# Studying magnetic circular vortex dichroism effect for photoionization of Rydberg atoms with vortex photons

Abstract
Rydberg atoms, renowned for their exceptional quantum properties, hold significant importance in quantum physics. The photoionization of Rydberg atoms serves as a critical tool for probing their unique characteristics. In this work, we investigate the photoionization dynamics of Rydberg atoms interacting with vortex photons—a class of structured light carrying intrinsic orbital angular momentum. This interaction gives rise to novel quantum phenomena distinct from conventional photoionization processes. Our results reveal that vortex photons exhibit exceptional sensitivity to the magnetic moments of Rydberg atoms, positioning them as a powerful spectroscopic tool for investigating Rydberg magnetism. It is also demonstrated that the initial photon energy must be carefully selected to observe significant experimental results. Furthermore, the photoionization process displays strong angular momentum selectivity, preferentially favoring configurations where the photon total angular momentum and atomic magnetic moment are aligned. This pronounced asymmetry directly manifests the chiral nature of the vortex photon-Rydberg atom interaction.
I Introduction
Since vortex photons were first shown to possess intrinsic orbital angular momentum (OAM) by Allen et al. in 1992 [1], they have found diverse applications across multiple domains. For instance, they serve as optical tweezers for nanoparticle manipulation [2, 3, 4], offer advantages in optical imaging compared to conventional methods [5, 6, 7, 8], enable high-capacity optical communication due to their large discrete degrees of freedom [9, 10, 11, 12], and provide ideal systems for studying quantum entanglement [13, 14, 15, 16], among other applications.
Optical activity represents another important field where vortex photons demonstrate significant value. The circular dichroism (CD) effect, which describes the different absorption or scattering of left- and right-handed circularly polarized light, can be generalized to vortex photons [17]. This leads to the so-called vortex dichroism (VD) effect, characterizing the distinct interaction of vortex photons with opposite OAM states with matter, as demonstrated both theoretically and experimentally [18, 19, 20]. When spin-orbit coupling is considered, the circular vortex dichroism (CVD) effect has also attracted considerable attention [21, 22, 23].
Magnetic dichroism effects, which put materials in magnetic fields and is the magnetic-field-induced counterparts of conventional dichroism effects, provide powerful tools for investigating magnetic properties. While magnetic circular dichroism (MCD)—characterizing the different absorption of left- and right-handed circularly polarized light in magnetic fields—is well established [24, 25, 26], magnetic vortex dichroism (MVD) which probes the dichroic response to vortex beams carrying opposite orbital angular momentum exhibits significantly stronger signals than MCD [27, 28, 29].
In this work, we investigate the magnetic circular vortex dichroism (MCVD) effect by comparing the response of polarized Rydberg atoms to vortex photon beams with opposite topological charges and helicities. Through parity analysis, this comparison is shown to be equivalent to reversing the atoms’ magnetic quantum numbers while keeping the photon beams unchanged. We theoretically calculate the photoionization cross-sections of Rydberg atoms interacting with vortex photons and present MCVD results for various initial topological charges and helicities. Specifically, initial vortex photon beam with energy $100$ eV is selected according to experimental condition [30, 31, 32]. The observed dependence of MCVD on the magnetic quantum number demonstrates the sensitivity of vortex photons to the magnetic moment of Rydberg atoms. Furthermore, asymptotic behavior of MCVD for vortex photon with large total angular momentum (TAM) shows angular momentum selectivity of the interaction.
The paper is organized as follows:
Section II presents the transition amplitude and cross-section calculations.
Section III provides MCVD results and analyzes the observed phenomena.
Section IV concludes the study.
Throughout this paper, three-dimensional vectors are denoted in boldface. $\hbar=c=1$ is used.
II Theoretical calculations
II.1 Transition matrix for Rydberg photoionization with initial plane wave photon
Under the first Born approximation, the single-electron photoionization matrix element for an atomic state with quantum numbers $(n,l,m_{l})$ by a plane wave photon in coordinate space is given by:
[…]
where:
  • $e$ is the electron charge
  • $m_{e}$ is the electron mass
  • $\mathbf{r}$ is the position coordinate
  • $\mathbf{k}$ is the incident photon momentum
  • $\mathbf{\epsilon}_{\mathbf{k}{\Lambda_{\gamma}}}$ is the photon polarization vector (helicity $\Lambda_{\gamma}$})
  • $\mathbf{p}$ is the outgoing electron momentum
  • $\hat{\mathbf{p}}=-i\mathbf{\nabla}$ is the momentum operator
  • $\psi_{f}$ is the final electron state
  • $\psi_{i}^{nlm_{l}}$ is the initial bound state
In spherical coordinates, the momenta are parameterized as $\mathbf{k}=(k,\theta_{\mathbf{k}},\phi_{\mathbf{k}})$ and $\mathbf{p}=(p,\theta_{\mathbf{p}},\phi_{\mathbf{p}})$. The photon polarization vector in Coulomb gauge takes the explicit form:
$\mathbf{\epsilon}_{\mathbf{k}{\Lambda_{\gamma}}}=\begin{pmatrix}-\dfrac{{\Lambda_{\gamma}}}{\sqrt{2}}e^{-i{\Lambda_{\gamma}}\phi_{\mathbf{k}}}\cos^{2}\dfrac{\theta_{\mathbf{k}}}{2}+\dfrac{{\Lambda_{\gamma}}}{\sqrt{2}}e^{i{\Lambda_{\gamma}}\phi_{\mathbf{k}}}\sin^{2}\dfrac{\theta_{\mathbf{k}}}{2}\\[10.0pt] -\dfrac{i}{\sqrt{2}}e^{-i{\Lambda_{\gamma}}\phi_{\mathbf{k}}}\cos^{2}\dfrac{\theta_{\mathbf{k}}}{2}-\dfrac{i}{\sqrt{2}}e^{i{\Lambda_{\gamma}}\phi_{\mathbf{k}}}\sin^{2}\dfrac{\theta_{\mathbf{k}}}{2}\\[10.0pt] \dfrac{{\Lambda_{\gamma}}}{\sqrt{2}}\sin\theta_{\mathbf{k}}\end{pmatrix}$. (2)
Choosing the final state as a plane wave $\psi_{f}^{*}=e^{-i\mathbf{p}\cdot\mathbf{r}}$, we obtain the momentum-space representation:
${\cal M}_{fi}^{P}=-\frac{e}{m_{e}}(\mathbf{\epsilon}_{\mathbf{k}{\Lambda_{\gamma}}}\cdot\mathbf{q})\tilde{\psi}_{i}^{nlm}(\mathbf{q}),$ (3)
where $\mathbf{q}\equiv\mathbf{p}-\mathbf{k}$ is the momentum transfer and $\tilde{\psi}_{i}$ denotes the Fourier transform of the initial state.
The Coulomb gauge condition $\mathbf{\epsilon}_{\mathbf{k}{\Lambda_{\gamma}}}\cdot\mathbf{k}=0$ implies:
$\mathbf{\epsilon}_{\mathbf{k}{\Lambda_{\gamma}}}\cdot\mathbf{q}$ $=\mathbf{\epsilon}_{\mathbf{k}{\Lambda_{\gamma}}}\cdot\mathbf{p}$
$=\frac{{\Lambda_{\gamma}}}{\sqrt{2}}p\Bigg{[}\sin\theta_{\mathbf{p}}\bigg{(}\sin^{2}\frac{\theta_{\mathbf{k}}}{2}e^{i{\Lambda_{\gamma}}(\phi_{\mathbf{k}}-\phi_{\mathbf{p}})}$
$\quad-\cos^{2}\frac{\theta_{\mathbf{k}}}{2}e^{-i{\Lambda_{\gamma}}(\phi_{\mathbf{k}}-\phi_{\mathbf{p}})}\bigg{)}+\sin\theta_{\mathbf{k}}\cos\theta_{\mathbf{p}}\Bigg{]},$ (4)
where $p\equiv\norm{\mathbf{p}}$ is the electron momentum magnitude.
The atomic wavefunction in momentum space has the closed-form expression [33]:
$\tilde{\psi}_{nlm_{l}}(\mathbf{q})$ $=2^{2l+2}l!(-i)^{l}n^{-l-2}\sqrt{\frac{2(n-l-1)!}{\pi Z^{3}(n+l)!}}\frac{Z^{2l+4}}{(q^{2}a^{2}+\delta^{2})^{l+2}}$
$\quad\times C_{n-l-1}^{l+1}\left(\frac{q^{2}a^{2}-\delta^{2}}{q^{2}a^{2}+\delta^{2}}\right)\left(\frac{qa}{Z}\right)^{l}Y_{l}^{m_{l}}(\theta_{\mathbf{q}},\phi_{\mathbf{q}}),$ (5)
with:
  • $Z$: nuclear charge
  • $a$: Bohr radius
  • $\delta\equiv Z/n$
  • $C_{n}^{\alpha}(x)$: Gegenbauer polynomial
  • $Y_{l}^{m_{l}}$: spherical harmonic function
II.2 Transition matrix for Rydberg photoionization with initial vortex photon
To demonstrate our key physical results, we focus specifically on alkali Rydberg atoms with the principal quantum number $n=l+1$. Single electron ionization is considered and thus $Z$ is set to be one.
II.2.1 Momentum Space Wavefunction for the selected Rydberg atoms
The atomic wavefunction in momentum space is given by:
$\tilde{\psi}_{nl{m_{l}}}(\mathbf{q})$ $=2^{2n}l!(-i)^{l}n^{-l-2}\sqrt{\frac{2}{\pi(2l+1)!}}\frac{q^{l}a^{l}Y_{l}^{m_{l}}(\theta_{\mathbf{q}},\phi_{\mathbf{q}})}{(q^{2}a^{2}+\delta^{2})^{l+2}}$
$=\hat{N}\frac{q^{l}Y_{l}^{m_{l}}(\theta_{\mathbf{q}},\phi_{\mathbf{q}})}{(n^{2}q^{2}a^{2}+1)^{l+2}},$ (6)
where the normalization constant is:
$\hat{N}=2^{2n}l!(-i)^{l}n^{l+2}a^{l}\sqrt{\frac{2}{\pi(2l+1)!}},$ (7)
and $\delta=1/n$ for Rydberg alkali atoms.
II.2.2 Plane Wave Photoionization for the selected Rydberg atoms
The transition amplitude for plane wave photons is:
${\cal M}_{fi}^{P}(\mathbf{k},\mathbf{p})$ $=-\frac{\hat{N}e{\Lambda_{\gamma}}}{\sqrt{2}m_{e}}\Bigg{[}p_{\perp}\sin^{2}\frac{\theta_{\mathbf{k}}}{2}{\rm e}^{i{\Lambda_{\gamma}}\phi}-p_{\perp}\cos^{2}\frac{\theta_{\mathbf{k}}}{2}{\rm e}^{-i{\Lambda_{\gamma}}\phi}$
$\quad+p_{z}\sin\theta_{\mathbf{k}}\Bigg{]}\frac{q^{l}Y_{l}^{m_{l}}(\theta_{\mathbf{q}},\phi_{\mathbf{q}})}{(n^{2}a^{2}q^{2}+1)^{l+2}},$ (8)
where $\phi=\phi_{\mathbf{k}}-\phi_{\mathbf{p}}$, $p_{\perp}=p\sin\theta_{\mathbf{p}}$ and $p_{z}=p\cos\theta_{\mathbf{p}}$. Specially, when photon momentum is parallel with polarization axis of the target atom (i.e. $\theta_{\mathbf{k}}=0$), we have
${\cal M}_{fi}^{P}(\mathbf{k},\mathbf{p})$ $=\frac{\hat{N}e{\Lambda_{\gamma}}}{\sqrt{2}m_{e}}p_{\perp}\frac{q^{l}Y_{l}^{m_{l}}(\theta_{\mathbf{q}},\phi_{\mathbf{q}})}{(n^{2}a^{2}q^{2}+1)^{l+2}}.$ (9)
This equation shows that $|{\cal M}_{fi}^{P}|$ is independent of photon helicity $\Lambda_{\gamma}$} under the first Born approximation.
For simplicity, we write Eq.II.2.2 as:
${\cal M}_{fi}^{P}(\mathbf{k},\mathbf{p})=N_{P}\mathcal{A}(\theta_{\mathbf{k}},\phi)\frac{q^{l}Y_{l}^{m_{l}}(\theta_{\mathbf{q}},\phi_{\mathbf{q}})}{(n^{2}a^{2}q^{2}+1)^{l+2}},$
where
$N_{P}=-\frac{\hat{N}e{\Lambda_{\gamma}}}{\sqrt{2}m_{e}}$
and
$\mathcal{A}(\theta_{\mathbf{k}},\phi)=p_{\perp}\sin^{2}\frac{\theta_{\mathbf{k}}}{2}{\rm e}^{i{\Lambda_{\gamma}}\phi}-p_{\perp}\cos^{2}\frac{\theta_{\mathbf{k}}}{2}{\rm e}^{-i{\Lambda_{\gamma}}\phi}+p_{z}\sin\theta_{\mathbf{k}}.$
II.2.3 Vortex Photon Photoionization for the selected Rydberg atoms
The most convenient vortex states are the Bessel vortex states which will be used in this paper. A Bessel vortex state for photon is
$\mathbf{A}^{B}(\mathbf{r},\,t)$ $=$ $\int\frac{d^{2}\vec{k}_{\perp}}{(2\pi)^{2}}\sqrt{\frac{2\pi}{\kappa}}(-i)^{m_{\gamma}}{\mathbf{\epsilon}}_{\mathbf{k}{\Lambda_{\gamma}}}\delta(|\vec{k}_{\perp}|-\kappa){\rm e}^{i\,({\mathbf{k}}\cdot{\mathbf{r}}-\omega t)}$ (10)
$=$ $N_{B}\int\frac{d\phi_{\mathbf{k}}}{2\pi}{\rm e}^{i\,m_{\gamma}\phi_{\mathbf{k}}}{\rm e}^{i\,\mathbf{k}\cdot\mathbf{r}}{\mathbf{\epsilon}}_{\mathbf{k}{\Lambda_{\gamma}}},$
where $N_{B}=(-i)^{m_{\gamma}}\sqrt{{\kappa}/{2\pi}}$, $\vec{k}_{\perp}$ is transverse part of momentum $\mathbf{k}$ (longitudinal part is written as $k_{z}$). In spherical coordinates, the momenta are parameterized as $\mathbf{k}=(k,\theta_{\mathbf{k}},\phi_{\mathbf{k}})$. There are four quantum numbers: energy $\omega$, transverse momentum $\kappa$, total angular momentum (TAM) quantum number $m_{\gamma}$ and helicity $\Lambda_{\gamma}$}. The state has TAM $\hbar m_{\gamma}$.
The position-dependent photon current density is [34] :
$j^{V}(r_{\perp})$ $=\frac{k\cos\theta_{\mathbf{k}}}{2\pi}\Bigg{[}\cos^{4}\frac{\theta_{\mathbf{k}}}{2}J_{m_{\gamma}-{\Lambda_{\gamma}}}^{2}(k_{\perp}r_{\perp})$
$\quad+\frac{1}{2}\sin^{2}\theta_{\mathbf{k}}J_{m_{\gamma}}^{2}(k_{\perp}r_{\perp})+\sin^{4}\frac{\theta_{\mathbf{k}}}{2}J^{2}_{m_{\gamma}+{\Lambda_{\gamma}}}(k_{\perp}r_{\perp})\Bigg{]},$ (11)
where $r_{\perp}$ denotes the transverse coordinate and $J_{n}(x)$ is the first kind Bessel function.
For central collision between vortex photon and Rydberg atom , the amplitude is obtained through azimuthal averaging:
${\cal M}^{V}_{fi}$ $=N_{B}\int_{0}^{2\pi}\frac{d\phi_{\mathbf{k}}}{2\pi}{\rm e}^{im_{\gamma}\phi_{\mathbf{k}}}M^{P}_{fi}(\mathbf{k},\mathbf{p})$
$=N_{B}N_{P}{\rm e}^{im_{\gamma}\phi_{\mathbf{p}}}\int_{0}^{2\pi}\frac{d\phi}{2\pi}{\rm e}^{im_{\gamma}\phi}\mathcal{A}(\theta_{\mathbf{k}},\phi)$
$\quad\times\frac{q^{l}Y_{l}^{m_{l}}(\theta_{\mathbf{q}},\phi_{\mathbf{q}})}{(n^{2}a^{2}q^{2}+1)^{l+2}}.$ (12)
Using the solid harmonic expansion:
$q^{l}Y_{l}^{m_{l}}(\theta_{\mathbf{q}},\phi_{\mathbf{q}})=\sqrt{\frac{2l+1}{4\pi}}R_{l}^{m_{l}}(\mathbf{q}),$ (13)
and the addition theorem for solid harmonics:
$R_{l}^{m_{l}}(\mathbf{r}+\mathbf{a})$ $=\sum_{\sigma=0}^{l}\sum_{\mu=-\sigma}^{\sigma}{\sqrt{\binom{2l}{2\sigma}}}R_{\sigma}^{\mu}(\mathbf{r})R_{l-\sigma}^{{m_{l}}-\mu}(\mathbf{a})$
[…]
where $\binom{2l}{2\sigma}$} is the binomial coefficient, we obtain the expanded formula:
$\frac{q^{l}Y_{l}^{m_{l}}(\theta_{\mathbf{q}},\phi_{\mathbf{q}})}{(n^{2}a^{2}q^{2}+1)^{l+2}}$ $=\frac{{\rm e}^{i{m_{l}}\phi_{\mathbf{p}}}}{(n^{2}a^{2}q^{2}+1)^{l+2}}$
$\quad\times\sum_{\sigma=0}^{l}\sum_{\mu=\mu_{\rm min}}^{\mu_{\rm max}}f_{l{m_{l}},\sigma\mu}(\mathbf{p})e^{i({m_{l}}-\mu)\phi},$ (15)
where the coefficient function is:
$f_{l{m_{l}},\sigma\mu}(\mathbf{p})$ $=\sqrt{\frac{(2l+1)(\sigma-\mu)!(l-\sigma-{m_{l}}+\mu)!}{4\pi(\sigma+\mu)!(l-\sigma+{m_{l}}-\mu)!}\binom{2l}{2\sigma}}$
$\quad\times p^{\sigma}(-k)^{l-\sigma}P_{\sigma}^{\mu}(\cos\theta_{\mathbf{p}})P_{l-\sigma}^{{m_{l}}-\mu}(\cos\theta_{\mathbf{k}})$
[…]
$P_{\sigma}^{\mu}$ is the associated Laguerre polynomial and the limitation of $\mu$ is: $\mu_{\rm min}={\rm Max}(-l+{m_{l}}+\sigma,\,-\sigma)$ and $\mu_{\rm max}={\rm Min}(l+{m_{l}}-\sigma,\,\sigma)$.
After substituting Eq.II.2.3 into Eq.II.2.3 and performing the variable change $z=e^{i\phi_{\mathbf{k}}}$, we obtain:
${\cal M}_{fi}^{V}$ $=\frac{i(-1)^{n}N_{B}N_{P}}{2\pi(na)^{2l+4}}e^{i({m_{l}}+m_{\gamma})\phi_{\mathbf{p}}}\sum_{\sigma=0}^{l}\sum_{\mu=\mu_{\rm min}}^{\mu_{\rm max}}f_{l{m_{l}},\sigma\mu}(\mathbf{p})$
$\quad\times\oint_{\mathcal{C}}dz\,\frac{\mathcal{B}(z)}{(p_{\perp}k_{\perp})^{l+2}z^{-\zeta}(z-z_{1})^{l+2}(z-z_{2})^{l+2}},$ (17)
where $\zeta={m_{l}}+m_{\gamma}+l-\mu$ and
$\mathcal{B}(z)=p_{\perp}\sin^{2}(\theta_{\mathbf{k}}/2)z^{1+{\Lambda_{\gamma}}}-p_{\perp}\cos^{2}(\theta_{\mathbf{k}}/2)z^{1-{\Lambda_{\gamma}}}+p_{z}\sin\theta_{\mathbf{k}}\,z,$
and $\mathcal{C}$ denotes the unit circle in the complex plane.
The contour integral in Eq. (II.2.3) can be evaluated using the residue theorem. The explicit analytical result is presented in Appendix A.
II.3 Differential Cross Section
The photoionization differential cross section is given by
$d\sigma=\frac{|{\cal M}|^{2}}{j}(2\pi)\delta(E_{f}-E_{i}-\omega)\frac{d^{3}\mathbf{p}}{(2\pi)^{3}},$ (18)
where $E_{f}$ is the kinetic energy of the final electron, $E_{i}$ is the binding energy of the initial electron, $\omega$ is the energy of the incident photon, and $j$ represents the incident photon current. For different initial photon state, current $j$ will be different.
[…]
$\frac{d\sigma}{d\theta_{\mathbf{p}}}=\frac{m_{e}p\sin\theta_{\mathbf{p}}}{2\pi j}|{\cal M}|^{2}.$ (19)
Here, for initial vortex photon, we use an effective photon current:
$j^{V}_{\rm eff}=\int j^{V}(r_{\perp})|\psi_{i}|^{2}(\mathbf{r})d^{3}\mathbf{r}.$
The efficiency of this current is discussed in Appendix.B.
III Results and Discussions
III.1 Magnetic circular-vortex dichroism effect
The magnetic circular dichroism (MCD) effect manifests distinct absorption or scattering probabilities for left- and right-circularly polarized photons incident on chiral targets in a static magnetic field. This phenomenon has emerged as a powerful tool for probing the magnetic properties of materials. For vortex photons, the analogous magnetic vortex dichroism (MVD) effect arises from different absorption or scattering of optical vortices carrying opposite orbital angular momentum (or topological charge). We introduce a novel dichroic phenomenon, termed magnetic circular-vortex dichroism (MCVD), which characterizes the different response of a magnetized target to polarized vortex beams with opposite TAM and helicity:
$\mathrm{MCVD}=\frac{\sigma(m_{\gamma},1)-\sigma(-m_{\gamma},-1)}{\sigma(m_{\gamma},1)+\sigma(-m_{r},-1)}$}.
Symmetry analysis
[…]
The underlying symmetry corresponds to a mirror inversion in the transverse plane, equivalent to a one-dimensional parity transformation. Under this operation, factors in Eq. II.2.3:
  • The phase factors ${\rm e}^{i\phi}/{\rm e}^{i\phi_{\mathbf{q}}}$ transform as ${\rm e}^{-i\phi}/{\rm e}^{-i\phi_{\mathbf{q}}}$ (up to a constant phase)
  • The length $q$ remains unchanged
  • The differential $d\phi$ changes sign ($d\phi\to-d\phi$)
[…]
$\mathrm{MCVD}(m_{\gamma},m_{l})$ $=\frac{\sigma(m_{\gamma},1,{m_{l}})-\sigma(-m_{\gamma},-1,{m_{l}})}{\sigma(m_{\gamma},1,{m_{l}})+\sigma(-m_{\gamma},-1,{m_{l}})}$
$=\frac{\sigma(m_{\gamma},1,{m_{l}})-\sigma(m_{\gamma},1,-{m_{l}})}{\sigma(m_{\gamma},1,{m_{l}})+\sigma(m_{\gamma},1,-{m_{l}})}.$ (20)
This demonstrates that MCVD measurement requires only target inversion along the photon’s propagation direction for experimental verification.
III.2 $m$ dependence of MCVD for Rydberg Atoms
To investigate the sensitivity of vortex photons to the magnetic moment of Rydberg atoms, we analyze the dependence of MCVD on the magnetic quantum number $m_{l}$} under two distinct experimental configurations. In both cases, we consider several types of vortex photons characterized by different TAM quantum number $m_{\gamma}$ and helicity $\Lambda_{\gamma}$}.
Figure 1: MCVD in Rydberg atom photoionization with vortex photons. The principal and orbital quantum numbers are fixed at $n=l+1=51$, while the magnetic quantum number varies from $10$ to $50$. Initial photon energy is chosen to be $100$ eV. Conical angle of vortex photon is $\theta_{k}=\pi/10$. Different lines corresponds to different cases with $m_{\gamma}$ ranging from 0 to $\pm 3$.
First Configuration: Fixed Principal and Orbital Quantum Numbers
In the first scenario, we fix the principal ($n$) and orbital ($l$) quantum numbers while varying the magnetic quantum number $m_{l}$}. Figure 1 presents the MCVD curves for different initial photon states, considering Rydberg atoms with $n=l+1=51$. For MCVD which we have defined in Eq. III.1, we examine vortex photons with TAM quantum number ranging from $m_{\gamma}=0$ to $m_{\gamma}=\pm 3$.
The MCVD approaches zero for the case $m_{\gamma}=1$, closely resembling the plane-wave photon scenario (Eq. 9) where the transition amplitude modulus becomes helicity-independent at tree level and the MCVD is zero. Other cases exhibit significant MCVD sensitivity to $m_{l}$}, directly reflecting the atomic magnetic moment. For $|m_{\gamma}|\geq 2$, the MCVD shows monotonic $m_{l}$}-dependence, confirming its effectiveness for Rydberg atom magnetic moment measurements. At large $m_{l}$}, MCVD asymptotically approaches 1, revealing enhanced chiral asymmetry for states with higher magnetic quantum numbers. The $|m_{\gamma}|=2$ case demonstrates stronger MCVD variations than $|m_{\gamma}|=3$, making it preferable for analysis. While cases with $m_{\gamma}=-1,\,0$ display non-monotonic trends, their high-$m_{l}$ tails show significant $m_{l}$ dependence.
Figure 2: MCVD in CRA photoionization with vortex photons. Targets follow $n=l+1={m_{l}}+1$ with $n$ varying from $11$ to $51$. Initial photon energy is chosen to be $100$ eV. Conical angle of vortex photon is $\theta_{k}=\pi/10$. Different lines corresponds to different cases with $m_{\gamma}$ ranging from 0 to $\pm 3$.
Second Configuration: CRAs With Variable Principal Quantum Numbers
In the second scenario, we investigate circular Rydberg atoms (CRAs) with varying principal quantum numbers $n$ while maintaining other parameters constant. With the magnetic quantum number fixed as ${m_{l}}=n-1$ (considering only ${m_{l}}\geq 0$), this configuration effectively studies $m_{l}$}-dependence. Figure 2 presents MCVD curves with $m_{l}$} ranging from $10$ to $50$. Again, for MCVD which we have defined in Eq. III.1, we examine vortex photons with TAM quantum number ranging from 0 to $\pm 3$.
The $m_{\gamma}=1$ case again shows MCVD approaching zero, demonstrating that vortex photons with TAM matching their helicity ($m_{\gamma}={\Lambda_{\gamma}}$) behave similarly to plane-wave photons. The $m_{\gamma}=-1$ and $m_{\gamma}=0$ cases exhibit both significant sensitivity to $m_{l}$} and monotonic dependence, confirming their utility for CRA magnetic moment measurements. For $|m_{\gamma}|=\pm 2$ and $\pm 3$ cases, MCVD saturation at 1 makes experimental identification challenging. This behavior contrasts with the case in the first configuration.
In conclusion, the MCVD exhibits a strong dependence on the magnetic moment of Rydberg atoms when interacting with vortex photons carrying specific, low values of TAM. For a general Rydberg atom configuration, the suitable TAM states are $m_{\gamma}=0,\,-1,\,\pm 2$. This selection is further restricted to $m_{\gamma}=0,\,-1$ for the CRA configuration. Experimental investigations should therefore focus primarily on these parameter regimes.
III.3 Photon-energy dependence of MCVD for Rydberg Atoms
Figure 3: MCVD in Rydberg atom photoionization with vortex photons. Targets follow $n=l+1=51$ and ${m_{l}}=20$. $E_{k}$ is energy of initial photon ranging from 10 eV to 100 eV. Conical angle of vortex photon is $\theta_{k}=\pi/10$. Different lines corresponds to different cases with $m_{\gamma}$ ranging from 0 to $\pm 3$.
In addition to the magnetic moment of Rydberg atoms, the MCVD is also influenced by the initial photon energy. Figures 3 and 4 demonstrate this dependence for the two configurations defined previously, using vortex photons with TAM quantum numbers $m_{\gamma}=0,\pm 1,\pm 2,\pm 3$.
For the first configuration (Fig. 3), with Rydberg atom quantum numbers $n=l+1=51$ and ${m_{l}}=20$, the MCVD modulus shows distinct behaviors: The cases $m_{\gamma}=-1$ and $m_{\gamma}=0$ exhibit significant enhancement with increasing photon energy, whereas the cases with $m_{\gamma}=1,\,\pm 2,\,\pm 3$ show invisible variation with photon energy varying.
Figure 4: MCVD in CRA photoionization with vortex photons. Targets follow $n=l+1={m_{l}}+1$ with $n=51$. $E_{k}$ is energy of initial photon ranging from 10 eV to 100 eV. Conical angle of vortex photon is $\theta_{k}=\pi/10$. Different lines corresponds to different cases with $m_{\gamma}$ ranging from 0 to $\pm 3$.
The second configuration (Fig. 4), with fixed principal quantum number $n=51$ for CRA, reveals similar trends: The cases $m_{\gamma}=-1$ and $m_{\gamma}=0$ again show strong energy dependence, while the other five cases remain largely unaffected. Notably, cases with $m_{\gamma}=-3$ and $m_{\gamma}=-2$ converge to 1-1 with overlapping curves, and cases with $m_{\gamma}=3$ and $m_{\gamma}=2$ similarly overlap while approaching $1$.
These results clearly indicate that the initial photon energy primarily affects MCVD for specific low values of TAM ($m_{\gamma}=0,-1$). Since these cases give the most pronounced magnetic moment dependence, careful selection of photon energy becomes crucial. This dual sensitivity underscores the importance of optimal parameter selection for experimental investigations of MCVD.
III.4 MCVD with different photon TAM $m_{\gamma}$
Figure 5: Dependence of MCVD on photon TAM quantum number $m_{\gamma}$ for different Rydberg atom configurations: (a) $n=l+1=51$ and $m_{l}=30$; (b) $n=l+1=51$ and $m_{l}=40$; (c) CRA with $n=31$; (d) CRA with $n=51$. All cases show $m_{\gamma}\in[-5,5]$. Initial photon energy is chosen to be $100$ eV. Conical angle of vortex photon is $\theta_{k}=\pi/10$.
To investigate the dependence of MCVD on the TAM quantum number $m_{\gamma}$ of the initial photon, we analyze fixed Rydberg atom states while varying $m_{\gamma}$. Fig. 5 presents four distinct cases:
  • Subfigures (a) and (b): Rydberg atoms with fixed $n=l+1=51$ at magnetic quantum numbers ${m_{l}}=30$ and ${m_{l}}=40$ respectively
  • Subfigures (c) and (d): CRAs (${m_{l}}=n-1$) with principal quantum numbers $n=31$ and $n=51$ respectively
All cases examine $m_{\gamma}$ ranging from 5-5 to $5$. Key observations include:
  • MCVD $\rightarrow 1$ when $m_{\gamma}\geq 2$
  • MCVD $\rightarrow-1$ when $m_{\gamma}\leq-2$
  • MCVD varies between 1-1 and $1$ for $1\leq m_{\gamma}\leq 1$
The asymtotic MCVD convergence to $\pm 1$ for $|m_{\gamma}|\geq 2$ is a common feature. This behavior reveals a fundamental chiral asymmetry in the photoionization dynamics of Rydberg atoms with vortex photons: At fixed TAM modulus, for Rydberg states with positive magnetic quantum number (${m_{l}}>0$), the ionization cross section is significantly enhanced when interacting with vortex photons carrying positive TAM quantum number ($m_{\gamma}\geq+2$) compared to those with negative ones ($m_{\gamma}\leq-2$). The parity-conserving interaction rules governing this process show a complementary preference for negative-$m$ states to couple more strongly with negative-$m_{\gamma}$ photons.
III.5 Discussions
We present a qualitative analysis of the observed phenomena, with circular Rydberg atoms (CRAs) serving as our primary example. While similar analysis can be extended to general Rydberg atoms, the latter case involves greater complexity.
III.5.1 Multiple expansion
We can perform a multiple expansion for the photon field, whereby Eq. (1) becomes
$\mathcal{M}_{fi}^{P}$ $=-N_{B}\frac{e}{m_{e}}\sum_{\sigma=0}^{\infty}\Bigg{[}\frac{1}{\sigma!}\int d^{3}\mathbf{r}\,e^{-i\mathbf{p}\cdot\mathbf{r}}(i\mathbf{k}\cdot\mathbf{r})^{\sigma}$
$\quad\times(\mathbf{\epsilon}_{k{\Lambda_{\gamma}}}\cdot\hat{\mathbf{p}})\psi_{i}(\mathbf{r})\Bigg{]}$
$=-N_{B}\frac{e}{m_{e}}\sum_{\sigma=0}^{\infty}\left[\frac{1}{\sigma!}(\mathbf{\epsilon}_{k{\Lambda_{\gamma}}}\cdot\mathbf{p})(-\mathbf{k}\cdot\mathbf{\nabla}_{\mathbf{p}})^{\sigma}\tilde{\psi}_{i}(\mathbf{p})\right],$ (21)
where
[…]
with $\phi=\phi_{\mathbf{k}}-\phi_{\mathbf{p}}$. In particular, the azimuthal dependence enters only in the form $\phi_{k}-\phi_{p}$ due to rotational invariance.
The transition matrix for the vortex case is
$\mathcal{M}^{V}_{fi}$ $=N_{B}\int\frac{d\phi_{\mathbf{k}}}{2\pi}e^{im_{\gamma}\phi_{\mathbf{k}}}\mathcal{M}^{P}_{fi}(\mathbf{k})$
$=-N_{B}\frac{e}{m_{e}}\sum_{\sigma=0}^{\infty}\Bigg{[}\frac{1}{\sigma!}\int\frac{d\phi}{2\pi}e^{im_{\gamma}\phi_{\mathbf{k}}}(\mathbf{\epsilon}_{k{\Lambda_{\gamma}}}\cdot\mathbf{p})$
$\quad\times(-\mathbf{k}\cdot\mathbf{\nabla}_{p})^{\sigma}\tilde{\psi}_{i}(\mathbf{p})\Bigg{]}$
$=-N_{B}\frac{e}{m_{e}}e^{im_{\gamma}\phi_{\mathbf{p}}}\sum_{\sigma=0}^{\infty}\Bigg{[}\frac{1}{\sigma!}\int\frac{d\phi}{2\pi}e^{im_{\gamma}\phi}(\mathbf{\epsilon}_{k{\Lambda_{\gamma}}}\cdot\mathbf{p})$
$\quad\times(-\mathbf{k}\cdot\mathbf{\nabla}_{\mathbf{p}})^{\sigma}\tilde{\psi}_{i}(\mathbf{p})\Bigg{]}.$ (23)
By considering Eq. (II.1), (III.5.1), and (III.5.1), we can analyze the dependence of multiple contributions on TAM of the vortex photon. For the $\sigma$-th order multiple contribution:
  • The factor $(\mathbf{\epsilon}_{\mathbf{k}{\Lambda_{\gamma}}}\cdot\mathbf{p})$ contains $\phi$-dependent phase terms $e^{im_{\Lambda}\phi}$ with $m_{\Lambda}=0,~\pm 1$
  • The factor $(-\mathbf{k}\cdot\mathbf{\nabla}_{p})^{\sigma}$ contains $\phi$-dependent phase terms with $m_{\nabla}=0,~\pm 1,~\cdots,~\pm\sigma$
The global $\phi$-dependent phase factor $e^{i(m_{\gamma}+m_{\Lambda}+m_{\nabla})\phi}$ serves as a filter when integrated over one complete cycle. This integration selects only terms satisfying the conservation condition: $m_{\gamma}+m_{\Lambda}+m_{\nabla}=0$, which constrains the photon TAM to $m_{\gamma}=-m_{\Lambda}-m_{\nabla}$, with allowed values spanning 0 to $\pm(\sigma+1)$. The multipole expansion exhibits the following characteristic hierarchy:
  • Dipole($\sigma=0$): $m_{\gamma}\in\{0,\pm 1\}$
  • Quadrupole($\sigma=1$): $m_{\gamma}\in\{0,\pm 1,\pm 2\}$
This pattern extends to higher multiple orders. For photo-ionization with vortex photon at fixed $m_{\gamma}$, the minimal contributing order is $\sigma=|m_{\gamma}|-1$ for $|m_{\gamma}|\geq 1$, while the $m_{\gamma}=0$ case begins with $\sigma=0$.
III.5.2 Approximations for CRAs
Since the operator $\partial_{p}$ always yields a $1/p$ term when acting on $\tilde{\psi}_{i}$, we can extract a global factor $k/p$ from Eq. III.5.1. This factor is typically small for non-relativistic photoionization processes. Consequently, higher-order terms generally has smaller contributions to Eq. III.5.1.
With the $\phi$ integral, we explicitly get $m_{\nabla}=-m_{\gamma}-m_{\Lambda}$. The lowest order contribution has $\sigma=|m_{\nabla}|=|m_{\gamma}+m_{\Lambda}|$. Eq. III.5.1 contains $\phi$-independent terms that has no contribution to the lowest order term, due to the filtering effect of $\phi$ integral. We can therefore neglect these terms when evaluating Eq. III.5.1, leading to the simplified expression:
[…]
For Rydberg atoms, the wave function in momentum space is given by Eq. 25:
$\tilde{\psi}_{nl{m_{l}}}(\mathbf{p})\propto\frac{p^{l}Y_{l}^{m_{l}}(\theta_{\mathbf{p}},\phi_{\mathbf{p}})}{(n^{2}p^{2}a^{2}+1)^{l+2}}.$ (25)
For Rydberg atoms with large $n$, the term $n^{2}p^{2}a^{2}$ is much larger than unity, allowing the approximation:
$\tilde{\psi}_{nl{m_{l}}}(\mathbf{p})\propto\frac{Y_{l}^{m_{l}}(\theta_{\mathbf{p}},\phi_{\mathbf{p}})}{n^{2}a^{2}p^{l+4}}.$ (26)
[…]
$-\mathbf{k}\cdot\mathbf{\nabla}_{\mathbf{p}}$ $\approx\frac{k\sin\theta_{\mathbf{k}}}{p}\left[(l+2)e^{-i\lambda\phi}+2e^{i\lambda\phi}\right],$
$(-\mathbf{k}\cdot\mathbf{\nabla}_{\mathbf{p}})^{2}$ $\approx\left(\frac{k\sin\theta_{\mathbf{k}}}{p}\right)^{2}\left[(l+2)e^{-i\lambda\phi}+2e^{i\lambda\phi}\right]$
$\quad\times\left[(l+2.5)e^{-i\lambda\phi}+2.5e^{i\lambda\phi}\right],$
$(-\mathbf{k}\cdot\mathbf{\nabla}_{\mathbf{p}})^{\sigma}$ $\approx\left(\frac{k\sin\theta_{\mathbf{k}}}{p}\right)^{\sigma}\prod_{j=1}^{\sigma}\Bigg{[}\left(l+\frac{j+3}{2}\right)e^{-i\lambda\phi}$
$\quad+\left(\frac{j+3}{2}\right)e^{i\lambda\phi}\Bigg{]}.$ (27)
The higher-order contributions are strongly suppressed by the $(k\sin\theta_{\mathbf{k}}/p)^{\sigma}$ factor, with the $\sin\theta_{\mathbf{k}}$ term providing additional suppression.
III.5.3 phenomenon analyzing
As established earlier, the $\phi$ integral in Eq. III.5.1 acts as a selection filter through the condition $m_{\gamma}+m_{\Lambda}+m_{\nabla}=0$, where $m_{\Lambda}=0$ or $\pm 1$ originates from the polarization term:
$\mathbf{\epsilon}_{\mathbf{k}{\Lambda_{\gamma}}}\cdot\mathbf{p}=p_{\perp}\sin^{2}\frac{\theta_{\mathbf{k}}}{2}e^{i{\Lambda_{\gamma}}\phi}-p_{\perp}\cos^{2}\frac{\theta_{\mathbf{k}}}{2}e^{-i{\Lambda_{\gamma}}\phi}+p_{z}\sin\theta_{\mathbf{k}}.$ (28)
The phase contributions $m_{\nabla}$ emerge from the operator $(-\mathbf{k}\cdot\mathbf{\nabla}_{\mathbf{p}})^{\sigma}$.
For $|m_{\gamma}|\geq 2$, the dipole contribution ($\sigma=0$) vanishes, and the dominant contribution arises from $\sigma=|m_{\nabla}|\geq 1$. According to Eq. III.5.2, four distinct cases emerge:
1. When $m_{\gamma}\geq 2$ (implying $m_{\gamma}+m_{\Lambda}>0$ and $m_{\nabla}<0$) and ${m_{l}}>0$ ($\lambda=1$):
$(-\mathbf{k}\cdot\mathbf{\nabla}_{\mathbf{p}})^{|m_{\nabla}|}$ $\rightarrow\left(\frac{k\sin\theta_{\mathbf{k}}}{p}\right)^{|m_{\nabla}|}\prod_{j=0}^{|m_{\nabla}|-1}\left(l+2+\frac{j}{2}\right)$
$\quad\times e^{-i|m_{\nabla}|\phi}$ (29)
2. When $m_{\gamma}\geq 2$ (implying $m_{\gamma}+m_{\Lambda}>0$ and $m_{\nabla}<0$) and ${m_{l}}<0$ ($\lambda=-1$):
$(-\mathbf{k}\cdot\mathbf{\nabla}_{\mathbf{p}})^{|m_{\nabla}|}$ $\rightarrow\left(\frac{k\sin\theta_{\mathbf{k}}}{p}\right)^{|m_{\nabla}|}\prod_{j=0}^{|m_{\nabla}|-1}\left(2+\frac{j}{2}\right)$
$\quad\times e^{-i|m_{\nabla}|\phi}$ (30)
3. When $m_{\gamma}\leq-2$ (implying $m_{\gamma}+m_{\Lambda}<0$ and $m_{\nabla}>0$) and ${m_{l}}<0$ ($\lambda=-1$):
$(-\mathbf{k}\cdot\mathbf{\nabla}_{\mathbf{p}})^{|m_{\nabla}|}$ $\rightarrow\left(\frac{k\sin\theta_{\mathbf{k}}}{p}\right)^{|m_{\nabla}|}\prod_{j=0}^{|m_{\nabla}|-1}\left(l+2+\frac{j}{2}\right)$
$\quad\times e^{i|m_{\nabla}|\phi}$ (31)
4. When $m_{\gamma}\leq-2$ (implying $m_{\gamma}+m_{\Lambda}<0$ and $m_{\nabla}>0$) and ${m_{l}}>0$ ($\lambda=1$):
$(-\mathbf{k}\cdot\mathbf{\nabla}_{\mathbf{p}})^{|m_{\nabla}|}$ $\rightarrow\left(\frac{k\sin\theta_{\mathbf{k}}}{p}\right)^{|m_{\nabla}|}\prod_{j=0}^{|m_{\nabla}|-1}\left(2+\frac{j}{2}\right)$
$\quad\times e^{i|m_{\nabla}|\phi}$ (32)
The significant difference in prefactors between these cases explains why CRAs with large $l$ exhibit strong directional preference: photoionization favors vortex photons co-rotating ($m_{\gamma}>0$ for ${m_{l}}>0$, $m_{\gamma}<0$ for ${m_{l}}<0$) with the atomic angular momentum. This manifests in MCVD values approaching $\pm 1$ (Figs. 2, Fig. 5). This reveals the chiral nature in photoionization of polarized Rydberg atoms using vortex photons.
For $|m_{\gamma}|=1$, the dipole contribution ($\sigma=0$) plays important role. Two scenarios occur:
1. When $m_{\gamma}$ and $\Lambda_{\gamma}$} share the same sign, the $p_{\perp}\cos^{2}(\theta_{\mathbf{k}}/2)e^{-i{\Lambda_{\gamma}}\phi}$ term dominates, i.e. dipole contribution dominates, yielding MCVD $\approx 0$ resembling initial plane wave case (see $m_{\gamma}=1$ case in Figs. 1 and 2).
2. When $m_{\gamma}$ and $\Lambda_{\gamma}$} have opposite signs, the $p_{\perp}\sin^{2}(\theta_{\mathbf{k}}/2)e^{i{\Lambda_{\gamma}}\phi}$ term, which gives dipole contribution, is suppressed by $\sin^{2}(\theta_{\mathbf{k}}/2)$. $\sigma=1,2$ contributions (given by $p_{z}\sin\theta_{k}$ and $p_{\perp}\cos^{2}(\theta_{\mathbf{k}}/2)e^{-i{\Lambda_{\gamma}}\phi}$ terms respectively) become important. Their interference with dipole contribution introduce significant $m_{l}$}-dependent MCVD (see $m_{\gamma}=-1$ case in Figs. 1 and 2).
For $m_{\gamma}=0$, the dipole contribution is given by $p_{z}\sin\theta_{k}$ term and suppressed by $\sin\theta_{k}$. The quadrupole contribution ($\sigma=1$) given by $p_{\perp}\cos^{2}(\theta_{\mathbf{k}}/2)e^{-i{\Lambda_{\gamma}}\phi}$ term can be comparable to the dipole contribution. Their interference introduce significant $m_{l}$}-dependent MCVD (see $m_{\gamma}=0$ case in Figs. 1 and 2).
IV Conclusions
Figures 1 and 2 demonstrate that the magnetic circular vortex dichroism (MCVD) is highly sensitive to the magnetic moment of Rydberg atoms when the incident vortex photons carry specific low values of total angular momentum (TAM). This sensitivity establishes optical vortices as a powerful tool for probing quantum magnetism in polarized Rydberg systems, paralleling their established role in studying optical magnetism in chiral materials [28]. Thus, vortex-based spectroscopy offers a novel approach for characterizing the magnetic properties of Rydberg atoms.
Experimentally, the MCVD of Rydberg atoms can be measured by irradiating the sample with identical vortex photon beams along the atomic polarization axis from opposite directions. The reversal of the vortex photons’ longitudinal momentum does not affect the measurement outcomes. Moreover, the energy of the initial vortex photon must be carefully selected to observe significant physical phenomena, as demonstrated in Figures 3 and 4.
We also show that he photoionization process exhibits strong angular momentum selectivity, preferentially favoring aligned configurations between photon TAM and atomic magnetic moment. The enhanced cross sections for co-rotating systems versus counter-rotating configurations provide definitive evidence of this angular momentum matching. This pronounced asymmetry directly reflects the chiral nature of vortex photon-Rydberg atom interactions.
Acknowledgments
This work was supported in part by Grants No. NSFC-12447117.
Appendix A Calculation of ${\cal M}^{V}$ with residue theorem
The contour integral in Eq. (II.2.3),
$\oint_{\mathcal{C}}\!dz\,\frac{\mathcal{B}(z)}{(p_{\perp}k_{\perp})^{l+2}z^{\mu-l-m_{\gamma}-{m_{l}}}(z-z_{1})^{l+2}(z-z_{2})^{l+2}},$ (33)
can be evaluated analytically via the residue theorem.
The integral has singularities at:
  • $z_{0}=0$ (inside unit circle)
  • $z_{1}=\frac{p^{2}+k^{2}-2p_{z}k_{z}+1/(n^{2}a^{2})-\sqrt{\Delta}}{2p_{\perp}k_{\perp}}$ (inside)
  • $z_{2}=\frac{p^{2}+k^{2}-2p_{z}k_{z}+1/(n^{2}a^{2})+\sqrt{\Delta}}{2p_{\perp}k_{\perp}}$ (outside)
where $\Delta=(p^{2}+k^{2}-2p_{z}k_{z}+1/(n^{2}a^{2}))^{2}-4p^{2}_{\perp}k^{2}_{\perp}$.
For the case that $\Lambda=1$, the transition matrix can be expressed as:
${\cal M}_{fi}^{V}$ $=\frac{i(-1)^{n}N_{B}N_{P}}{2\pi(na)^{2l+4}}e^{i({m_{l}}+m_{\gamma})\phi_{p}}$
$\quad\times\sum_{\sigma=0}^{l}\sum_{{\mu=\mu_{\rm min}}}^{{\mu_{\rm max}}}f_{l{m_{l}},\sigma\mu}(\mathbf{p})(X_{a}^{+}+X_{b}^{+}+X_{c}^{+}),$ (34)
where the residue contributions are:
$X^{+}_{a}$ $=\frac{p_{\perp}\sin^{2}\frac{\theta_{\mathbf{k}}}{2}}{(p_{\perp}k_{\perp})^{l+2}}\oint dz\frac{1}{z^{-2-\zeta}(z-z_{1})^{l+2}(z-z_{2})^{l+2}}$ (35)
$X^{+}_{b}$ $=\frac{p_{z}\sin\theta_{\mathbf{k}}}{(p_{\perp}k_{\perp})^{l+2}}\oint dz\frac{1}{z^{-1-\zeta}(z-z_{1})^{l+2}(z-z_{2})^{l+2}}$ (36)
$X^{+}_{c}$ $=\frac{-p_{\perp}\cos^{2}\frac{\theta_{\mathbf{k}}}{2}}{(p_{\perp}k_{\perp})^{l+2}}\oint dz\frac{1}{z^{-\zeta}(z-z_{1})^{l+2}(z-z_{2})^{l+2}},$ (37)
where $\zeta={m_{l}}+m_{\gamma}+l-\mu$ and $X^{+}_{a}$ has residues
$\mathrm{Res}(X^{+}_{a},z_{1})$ $=\frac{2\pi ip_{\perp}\sin^{2}\frac{\theta_{\mathbf{k}}}{2}\binom{-l-2}{l+1}}{(p_{\perp}k_{\perp})^{l+2}z_{1}^{-2-\zeta}(z_{1}-z_{2})^{2l+3}}$
$\quad\times{}_{2}H_{1}\Big{(}-2-\zeta,-l-1,-2l-2,1-\frac{z_{2}}{z_{1}}\Big{)},$ (38)
with $_{2}H_{1}(a,b,c,d)$ being the Confluent hypergeometric function and
$\mathrm{Res}(X^{+}_{a},z_{0})$ $=\frac{2\pi ip_{\perp}\sin^{2}\frac{\theta_{\mathbf{k}}}{2}\binom{-l-2}{-3-\zeta}}{(p_{\perp}k_{\perp})^{l+2}(-z_{1})^{l+2}(-z_{2})^{\mu-{m_{l}}+m_{\gamma}-1}}$
$\quad\times{}_{2}H_{1}\Big{(}l+2,3+\zeta,2+\zeta,\frac{z_{2}}{z_{1}}\Big{)}$
$\text{only for }m_{\gamma}\leq-3-l;$ (39)
$X^{+}_{b}$ has residues
$\mathrm{Res}(X^{+}_{b},z_{1})$ $=\frac{2\pi ip_{z}\sin\theta_{\mathbf{k}}\binom{-l-2}{l+1}}{(p_{\perp}k_{\perp})^{l+2}z_{1}^{-1-\zeta}(z_{1}-z_{2})^{2l+3}}$
$\quad\times{}_{2}H_{1}\Big{(}-1-\zeta,-l-1,-2l-2,1-\frac{z_{2}}{z_{1}}\Big{)}$ (40)
and
$\mathrm{Res}(X^{+}_{b},z_{0})$ $=\frac{2\pi ip_{z}\sin\theta_{\mathbf{k}}\binom{-l-2}{-2-\zeta}}{(p_{\perp}k_{\perp})^{l+2}(-z_{1})^{l+2}(-z_{2})^{\mu-{m_{l}}+m_{\gamma}}}$
$\quad\times{}_{2}H_{1}\Big{(}l+2,2+\zeta,1+\zeta,\frac{\delta_{2}}{\delta_{1}}\Big{)}$
$\text{only for }m_{\gamma}\leq-2-l;$ (41)
$X^{+}_{c}$ has residues
$\mathrm{Res}(X^{+}_{c},z_{1})$ $=\frac{-2\pi ip_{\perp}\cos^{2}\frac{\theta_{\mathbf{k}}}{2}\binom{l+1}{-l-2}}{(p_{\perp}k_{\perp})^{l+2}z_{1}^{-\zeta}(z_{1}-z_{2})^{2l+3}}$
$\quad\times{}_{2}H_{1}\Big{(}-\zeta,-l-1,-2l-2,1-\frac{z_{2}}{z_{1}}\Big{)}$ (42)
and
$\mathrm{Res}(X^{+}_{c},z_{0})$ $=\frac{-2\pi ip_{\perp}\cos^{2}\frac{\theta_{\mathbf{k}}}{2}\binom{-l-2}{-1-\zeta}}{(p_{\perp}k_{\perp})^{l+2}(-z_{1})^{l+2}(-z_{2})^{\mu-{m_{l}}+m_{\gamma}+1}}$
$\quad\times{}_{2}H_{1}\Big{(}l+2,1+\zeta,\zeta,\frac{z_{2}}{z_{1}}\Big{)}$
$\text{only for }m_{\gamma}\leq-1-l.$ (43)
For the case that ${\Lambda_{\gamma}}=-1$, we just make exchange $\sin^{2}(\theta_{\mathbf{k}}/2)\rightleftharpoons-\cos^{2}(\theta_{\mathbf{k}}/2)$ from ${\Lambda_{\gamma}}=1$ case.
Appendix B Photon current for vortex photon-Rydberg atom collisions
B.1 Photon Current in Literature
Reference [34] provides detailed expressions for photon current densities in vortex photon collisions with ground state atoms, covering three distinct cases: macroscopic targets, mesoscopic targets, and single-atom targets.
Plane Wave Case
The photon current density is spatially uniform:
$j^{P}=\frac{k}{2\pi}.$
This expression appears in the denominator of the cross section formula (Eq. 18).
Vortex Case – Macroscopic Target
Considering Bessel vortex states incident on a macroscopic target (large radius with uniformly distributed atoms), the position-dependent photon current density is:
$j^{V}(r_{\perp})$ $=\frac{k\cos\theta_{\mathbf{k}}}{2\pi}\Big{[}\cos^{4}(\theta_{\mathbf{k}}/2)J_{0}^{2}(k_{\perp}r_{\perp})$
$\quad+\frac{1}{2}\sin^{2}\theta_{\mathbf{k}}J_{1}^{2}(k_{\perp}r_{\perp})+\sin^{4}(\theta_{\mathbf{k}}/2)J^{2}_{2}(k_{\perp}r_{\perp})\Big{]},\$(44)
where $r_{\perp}$ denotes the transverse coordinate. In the limit $\theta_{\mathbf{k}}\rightarrow 0$, $j^{V}(r_{\perp})\rightarrow j^{P}$.
For cross section calculations, we use the coordinate-independent effective current:
$j^{V}_{\text{mac}}=\int j^{V}(b)d^{2}\vec{b}=\frac{Rk\cos\theta_{\mathbf{k}}}{\pi k_{\perp}},$
where $R$ is the target radius and $\vec{b}$ represents the impact parameter. The radius $R$ becomes irrelevant for sufficiently large targets.
Vortex Case – Mesoscopic and Single-Atom Targets
For mesoscopic targets (small radius with atomic distribution $n(\vec{r}_{\perp})$), the effective photon current becomes:
$j^{V}_{\text{mes}}=\int j^{V}(r_{\perp})n(\vec{r}_{\perp})d^{2}\vec{r}_{\perp}.$
For a single-atom target treated as a point particle (e.g., a ground-state atom), we take $n(\vec{r}_{\perp})=\delta(r_{\perp})$, yielding:
$j^{V}_{\text{sin}}=\int j^{V}(r_{\perp})\delta(r_{\perp})d^{2}\vec{r}_{\perp}=\frac{k\cos\theta_{\mathbf{k}}\cos^{4}(\theta_{\mathbf{k}}/2)}{2\pi}.$
B.2 Photon Current for Vortex Beams with Rydberg Atom Targets
Rydberg atoms, having large spatial extent, cannot be treated as point particles. Following the methodology from previous cases, we define the effective photon current as:
$j^{V}_{R}=\int j^{V}(r_{\perp})|\psi_{i}|^{2}(\vec{r})d^{3}\mathbf{r}.$
This formulation correctly reduces to:
  • Plane wave case (through wavefunction normalization)
  • Macroscopic case: $j^{V}_{\text{mac}}=\int j^{V}(r_{\perp})\delta(z)d^{3}\mathbf{r}$
  • Mesoscopic case: $j^{V}_{\text{mes}}=\int j^{V}(r_{\perp})n(\vec{r}_{\perp})\delta(z)d^{3}\mathbf{r}$
The $\delta(z)$ terms reflect the longitudinal localization of conventional atomic targets. Physically, the effective current represents the photon flux experienced by the target.